consider the hydrogen atom. we assume that the atom is spherical and that the proton is also spherical. let Dh and Vh be the density and volume of the hydrogen atom, respectively. also, let Dp and Vp be the density and the volume of the proton, respectively. show that mathematically: DhVh = DpVp
1
Expert's answer
2019-11-14T06:17:09-0500

Assuming that the hydrogen atom is spherical;

Density Dh=m/v

Volume Vh= m/Dh but v=Vh.

Therefore, mass m of hydrogen atom is given by;

m= Dh*Vh =(m/Vh) *(m/Dh)

For proton;

Density Dp= m/v

Volume Vp= m/DP but v=Vp.

Therefore, mass m of hydrogen proton is given by;

m=Dp*Vp =(m/VP) *(m/DP)

Hence; DhVh=DpVp